OpenStudy (anonymous):

Find the slope m of a line parallel to the line that passes through (1, 2) and (2, 3).

OpenStudy (the_fizicx99):

M = y2 -y1/x2-x1 Latex was a fail u.u

OpenStudy (anonymous):

use slope formula to find slope

OpenStudy (the_fizicx99):

Says the slope is 1 ._. So rise 1 run 1

OpenStudy (anonymous):

\[M=\frac{ 3-2}{ 2-1 } ?\]

OpenStudy (the_fizicx99):

yes

OpenStudy (anonymous):

then use point slope equation to find the equation

OpenStudy (the_fizicx99):

\[M = \frac{ y2 - y1 }{ x2 - x1 } = \frac{ 3-2 }{ 2-1 } = \frac{ 1 }{ 1 } = 1\]

OpenStudy (anonymous):

So, the answer will be M=1

OpenStudy (anonymous):

Thank you all so much ! :)